OpenStudy (unknownrandom):

Need help solving this inequality. (1/2p)+(3/4p)<(1/2)

OpenStudy (anonymous):

\[\dfrac{a}{b}+\dfrac{c}{d}=\dfrac{ad+bc}{bd}\] use this rule

OpenStudy (unknownrandom):

I am unfamiliar with that formula which letters are which? @Jonask

OpenStudy (anonymous):

\[\dfrac{1}{2p}+\dfrac{3}{4p}=\dfrac{6p}{8p}=\dfrac{3}{4}\le\dfrac{2}{4}=\dfrac{1}{2}\]

OpenStudy (unknownrandom):

Thanks is that supposed to be 2 1/2? @Jonask

OpenStudy (anonymous):

its so is it \[\dfrac{3}{4}> \dfrac{2}{4}\] since \(3> 2\) i think its supposed to be more than,not less than